During a certain thermodynamic process, 418 J of work are done on a system and 214 cal of heat are transferred to the system. The change in internal energy during the process is
Answer #1

first Caw of thermodynamics Do = q + w w=418 5 (on the system) q = 214 cal = 914x 2.184 J E I cal = 1.1845] AU = q + w = [418